Calculus II, Spring 2016 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 16

1. Section 16.5: Problems 5, 11, 13, 19, 31, 33, 49, 51, 55, 56.

2. Section 16.6: Problems 17, 19, 21, 25, 35, 37.